OpenStudy (anonymous):

What is the value of the expression on -4 x to the 3rd power when x=-2? I got -32 and when submitted its said it was wrong and they say 32 is right?

OpenStudy (anonymous):

can u explain this to me please @jim_thompson5910

jimthompson5910 (jim_thompson5910):

-4x^3 = -4*x*x*x -4x^3 = -4*(-2)*(-2)*(-2) .... replace x with -2 -4x^3 = 8*(-2)*(-2) -4x^3 = -16*(-2) -4x^3 = 32

OpenStudy (anonymous):

i wish i had your brain thank you again

jimthompson5910 (jim_thompson5910):

I recommend breaking it down like that until you get the hang of it

OpenStudy (anonymous):

yeah thats what im doing right now trying to figure it out